OpenStudy (anonymous):

Medal Here!!! Can you arrange four 9's and use of atmost 2 math symbols , make the total be 100? Good Luck.

OpenStudy (anonymous):

99+9/9

OpenStudy (anonymous):

Wrong.

OpenStudy (anonymous):

four 9's.....2 math symbols....and equal 100...looks right to me...

OpenStudy (anonymous):

99+9/9 = 12

OpenStudy (anonymous):

Please tell me ur joking......

OpenStudy (anonymous):

9/9=1 +99=100 just switch them bud 9/9+99

OpenStudy (anonymous):

Nope. Your answer is wrong.

OpenStudy (anonymous):

It should be the two 9's are collide to each other to be able to have four 9's arranged with 2 math symbols.

OpenStudy (anonymous):

Idk who you are, but that is the ONLY answer I see, and no that's wrong, I know im right, do u want proof? ill show it to you

OpenStudy (anonymous):

Sure. Then I'll show mine.

OpenStudy (anonymous):

http://solution-dailybrainteaser.blogspot.com/2012/09/tricky-equation-puzzle.html Scroll down and see the discussion.

OpenStudy (anonymous):

99/.99 =100

OpenStudy (anonymous):

Well that's the correct answer. 99/.99 not 9/9+99 maybe this can be also but it is easier to do the suggested answer.

OpenStudy (anonymous):

Alright, medal please man /_^

OpenStudy (anonymous):

can i get medaled i got it right